OpenStudy (anonymous):

Use substitution to solve. x-5y=10 2x-10y=20

jimthompson5910 (jim_thompson5910):

Hint: solve the first equation for x x-5y=10 x=10+5y then plug this into the second equation 2x-10y=20 2(10+5y)-10y=20 20+10y-10y=20 I'll let you finish

OpenStudy (anonymous):

They both cancel out so it would equal 0,0

OpenStudy (anonymous):

Do you know how to use elimination and solve this question? X+4y=-8 X-4y=-8